Simulation of Shielding Options for an Active Collimator System

Christopher Gauthier University of Connecticut

(Dated: March 13, 2003)

Abstract

This paper covers the details of a simulation of three shielding options for an active collimator system to be used in the upcoming GlueX experiment. The simulations were run with GEANT, a high-energy particle physics simulation program developed at CERN. Three collimator shielding options, heavy, moderate and light shielding, were simulated. The simulations revealed that the light shielding option was the least effective in blocking secondary particles from penetrating into the experimental hall while the heavy and moderate shielding options both provided adequate shielding for the collimator. The heavy and moderate shielding options lead to acceptable background levels in the detectors, however, cost and ease of construction make the moderate shielding option the more desirable of the three.

INTRODUCTION

The GlueX Experiment will test our current picture of the subatomic world. The force between quarks is thought to be caused by a nuclear field called ”glue”. Although the properties of the glue are well understood where the field is weak, inside the nucleus where this field is quite strong its properties are largely unknown. In the GlueX experiment, a beam of high energy gamma rays will excite this glue, thereby revealing its physical properties at high energies. The GlueX experiment will rely on the use of high energy gamma rays to probe the internal structure of the bond between quarks. Unpolarized gamma rays can scatter into many more final states than a linearly polarized beam, which would needlessly complicate a partial wave analysis of the scattered beam. An initial state of linearly polarized gamma rays limits the number of possible final states, which makes a partial wave analysis easier and reduces the number of statistics needed. Using a beam of linearly polarized gamma rays in the GlueX experiment would improve the precision of the results.

The Coherent Bremsstrahlung process creates a beam of gamma rays with a degree of linear polarization that is dependent on the angle of emission. This angular dependence can be used to select and remove the unpolarized component of the gamma ray beam. A collimator can absorb a significant amount of the unpolarized gamma rays, while leaving the polarized unaffected. Increasing the polarization of the gamma rays through collimation has the adverse effect of decreasing the total flux reaching the target. It has been determined that collimating the beam to 40% polarization will reduce the flux to 20%, with 80% of the beam’s total power absorbed by the collimator. The process of collimation has the disadvantage of creating copious amounts of unwanted secondary particles.

This experiment will use a charged particle track detector that is sensitive to low energy particles, therefore, effective shielding for the collimator is necessary to keep any background from interfering with the performance of this detector. Three different shielding scenarios for the collimator system have been designed. For each design, a simulated beam of coherent bremsstrahlung photons passed through the collimator and the flux of secondary particles reaching the tracker was measured. The results for each design were compared with each other and with the cosmic ray background levels to determine which option worked the best at shielding the tracker from secondary particels.

SOFTWARE

The simulation was carried out using Geant, a physics simulation program developed at CERN. Geant simulates the passage of high-energy particles through matter using the Monte Carlo technique. The program uses a random number sequence to assign each particle a unique track history based on known interaction probabilities. For example, as a photon traverses a given material, for each type of physical process that photons can undergo in that medium, a ”length until interaction” is calculated, with a probability of occurrence given by the exponential distribution whose average is determined by the cross section. The process with the shortest ”length until interaction” will be the process that takes place. A similar procedure determines the types and four momenta of the final particles produced by the interaction. The original Geant release includes many of the common processes that occur in high-energy particle physics. Some interactions that are important in this analysis are not included in the standard Geant release these include photo-hadronic interactions and muon pair production. Art Snyder from SLAC has created a Geant plugin package called Gelhad that simulates the photoproduction of neutrons, protons, alpha particles, and pions. To simulate the production of muon pairs by photons, the electron-positron pair production package Gpairg was modified to convert a fraction of the electron-positron pairs into muon pairs. The fraction of electron positron pairs lost by this algorithm is on the order of (m_{e}/m_{μ})^{2} = 2.34×10^{−5} , therefore the decrease in the electron positron background level is negligible. Seven virtual detector volumes named DET1 through DET7 are placed in the virtual GlueX environment in areas that are of the most interest. Each time a particle passes through any one of these virtual detectors the particle’s position, momentum, energy and species are recorded in one of seven ntuples, one ntuple for each virtual detector. This system cuts down on the amount of disk space that a completed simulation uses, and allows for quicker access of data. The placement of the virtual detectors in the simulated GlueX environment is shown in Fig. 1. The static properties of the GlueX experimental volume, which includes the design parameters for the collimator system and the detector assembly, are stored in HDDS, an XML-based geometry specification. In HDDS, the geometry of a

FIG. 1: Figure showing the placement of the virtual detectors in the simulated GlueX environment. Virtual detector placement is the same for all three shielding options

virtual object is declared by the user in an XML file by the use of XML tags. For example, the tag

         <box name="BOX1" X_Y_Z="10.0  10.0  10.0" material="Air"/>

creates a box in Geant named BOX1 with dimensions 10.0 cm by 10.0 cm by 10.0 cm and is filled with air. The tag

         <composition name="CompBox" envelope="MBOX">
            <posXYZ volume="BOX1" X_Y_Z="0.0  0.0  0.0" />
         </composition>

positions BOX1 at the origin of a mother volume MBOX and then declares the composite volume as CompBox. Each subsystem of GlueX has its own XML file description, and these are linked and converted into a FORTRAN file using a program, hdds-geant.

THREE SCENARIOS

The basic collimator system itself is composed of two separate stages. When the gamma rays reach the first (primary) collimator, rays that left the source at large enough angles are absorbed on the collimator face while allowing gamma rays with low emission angles to pass through unscathed. The polarization of gamma rays created through Coherent Bremsstrahlung depend sensitively on the emission angle; therefore highly polarized gamma rays will pass through the collimator, while unpolarized gamma rays will be absorbed on the collimator face. To reduce the background further, a sweeping magnet with a strength of 0.4 T·m is placed in front of the collimator along with a large iron absorber to stop charged particles. This process is repeated by a second stage in the collimator assembly. Simulations of three shielding scenarios were run; each incorporating the same basic twocollimator design. The shielding designs have been designated as heavy, moderate and light, to distinguish between their levels of shielding. The dimensions of the hall are the same as given in the most recent blueprints of HallD.

Heavy Shielding Scenario

The Heavy shielding option shown in Fig. 2, utilizes sufficient shielding to stop essentially all particles produced by collimation. The collimator assembly is encased by a thick concrete shield that is, at its thinnest, 50 cm and at its thickest, 100 cm. The iron absorbers in front of the first and second sweeping magnets are 5.0 m and 3.0 m long, respectively. A lead wall with dimensions 5.5×4.2×.15m is placed in front of the entire collimator assembly for additional shielding. The dimensions of the heavy shielded collimator components and their placement relative to the detector assembly are shown in Table I.

Moderate Shielding Scenario

The moderate shielding option shown in Fig. 3 is a variation of the heavy shielding scenario. The front of the concrete shield has been reduced to a thickness of 35 cm and the lead plate has been reduced in size to the dimensions 4.7×2.9×.15 m. To compensate for this loss in shielding , two large concrete blocks are placed in front of the first and second absorbers. Both the absorbers themselves are shorter in this shielding version. The first absorber is now 2.0 m in length and the second absorber is now 1.0 m in length. The design parameters for the moderate shielding option are shown in Table II.

Light Shielding Scenario

The light shielding option shown in Fig. 4 is a further reduction of the heavy shielding option. The lead plate is now 4.5×2.7×.15 m and is inside the concrete shield right in front of the second absorber. The first and second iron absorbers are 2.0 m and 1.0 m in length, respectively. Similar to the moderate shielding design, the thickness of the concrete shield is thinner than in the heavy shielding option, here the concrete in the front of the collimator cave is 35 cm. The design parameters for this scenario are shown in Table III.

RESULTS

Simulations of each of these designs were run starting with 30 million original photons produced through a coherent bremsstrahlung generator, this corresponds to 7.0 ms of run time at 1.0 μA. The simulation was used to measure the flux as a function of the radial distance from the beam line for photons, electron-positron pairs, neutrons, and muon pairs at DET7. The charged particle tracker, shown in Fig. 5, is sensitive to low energy particles, therefore, an analysis of the background levels of electrons and muons will be an important factor in determining the best shielding option. Only particles with a radial distance greater than 5 cm were included in this analysis since photons at large angles are the ones that the barrel calorimeter must be shielded from and because it is important to know how much background is reaching the particle tracker, which has an inner radius of 5 cm centered on the beamline. The fraction of counts made by the tracker due to background is given by (1e^{−B·δt}), where B is the average number of background particles entering the tracker per unit time and δt is the shutter time of the tracker, which is roughly equal to 150 ns. After several simulations, it was discovered that the background could be significantly reduced with the placement of a vacuum pipe that extends from the face of the second sweeping magnet to the face of the GlueX detector assembly. The vacuum pipe in each simulation was made of an iron tube 3 mm thick and was closed at both ends by a plastic Mylar cap that was 0.5 mm thick. Study of different arrangements of the collimator components showed that the background flux depends on the separation between the primary and secondary collimators. This is caused by the decrease in the distance between the secondary collimator and the detector assembly, which increases the intensity of the particle shower since the intensity decreases with increasing distance from its source. Analysis of the background flux lead to the conclusion that much of the background is emitted at small angles, therefore much emphasis was put on shielding the front of the collimator cave.

Heavy Shielding

Figure 6 and Table IV show the results of the heavy shielding option. The heavy shielding scenario had the lowest total flux of all the shielding options. The rate of secondary particles entering the tracker was approximately ten thousand times greater than the rate of cosmic rays entering the tracker. The muon flux was very small relative to the electron-positron flux and was not a significant component of the total background reaching the tracker. The flux as a function of radius decreases quickly as the radial distance increases. This indicates that there is sufficient shielding in the front of the collimator cave.

Moderate Shielding

The background flux in the moderate shielding option, as shown in Fig. 7 and Table V, was very close to the levels in the heavy shielding option. There were only slight increases from the heavy shielding option in the background levels of the electrons, photons and muons and a decrease in the neutron flux. The decrease in neutron flux is a result of the two concrete blocks in front of the iron absorbers, which absorb excess neutrons. From the background rates at the tracker given in Tables 4 and 5, it is clear that the moderate shielding option is better at shielding the tracker than the heavy shielding option, which is again a result of the two concrete blocks.

Light Shielding

As indicated in Fig. 8 and Table VI, there was a big increase in the flux levels for each particle type. Some particle fluxes increased by two orders of magnitude. This significant increase in flux is not only the result of a reduction in the shielding but also the increased distance between the primary and secondary collimators. The decrease in the flux as a function of distance is considerably slower in this scenario, which indicates a lack of shielding in the front of the collimator cave.

CONCLUSION

These simulated collimators do not take into account all aspects of a real active collimator system. The simulations do not take into account an access tunnel into the collimator housing, which could change the background levels upstream. However, an access tunnel would most likely be placed in the center of the collimator housing where only particles with large emission angles would escape, therefore no noticeable increase in the background flux would be seen at the detectors. Other components of the collimator system that were left out of the simulation were fixtures to hold the collimator components in place, and a flux monitor inside the primary collimator to help position the collimator on the beamline. These components were seen as irrelevant to this study since little of the initial gamma ray beam flux would be absorbed by these components. The photohadronic interaction generator Gelhad, which was used in this simulation, should be checked against other hadronic generators to confirm Gelhad’s reliability. Muon-nuclear interactions were not included in this study, since it is a negligible second order effect. The electromagnetic generator in Geant is well tested and has proven itself reliable. From this study, the light shielding option does not have enough shielding to stop a large amount of the background produced in the collimator. Therefore, the light shielding scenario would not be the most desirable option. The differences in the flux between the heavy and moderate shielding scenarios were negligible. The heavy shielding option, although having the lowest total flux, did not shield the tracker as effectivly as the moderate shielding option could. The moderate shielding scenario would be more cost effective and easier to manufacture than the heavy shielding option and this coupled with low background levels at the detector assembly make the moderate shielding option the better of the three scenarios.

FIG. 2: Diagram of the heavy shielded collimator FIG. 3: Diagram of the moderate shielded collimator FIG. 4: Diagram of the light shielded collimator FIG. 5: Diagram of the GlueX detector assembly

TABLE I: Dimensions of the heavy shielded collimator components.

Component  Material Width Height Length Distance from detector
(shape: cubic)  (cm)  (cm)  (cm)  assembly (cm)
concrete housing  concrete  550  420  1200  na
inner collimator cave  air  450  270  1050  na
first sweeping magnet  iron  20  30  50  2070
first absorber  iron  50  50  500  1560
second sweeping magnet  iron  20  30  50  1440
second absorber  iron  100  100  300  1130
lead wall  lead  550  420  15  980

Component  Material  r_{in}  r_{out}  Length Distance from detector
(shape: cylinderical)  (cm)  (cm)  (cm)  assembly (cm)
primary collimator  tungsten  0.25  10  30  2130
secondary collimator  nickel  0.50  10  50  1500
vacuum pipe  iron  2.2  2.5  1488  na

TABLE II: Dimensions of the moderate shielded collimator components.

Component  Material Width Height Length Distance from detector
(shape: cubic)  (cm)  (cm)  (cm)  assembly (cm)
concrete housing  concrete  550  420  1200  na
inner collimator cave  air  450  270  1150  na
first sweeping magnet  iron  20  30  50  2020
first absorber  iron  50  50  200  1810
first concrete block  concrete  200  200  100  1670
second sweeping magnet  iron  20  30  50  1390
second absorber  iron  100  100  100  1280
second concrete block  concrete  430  250  100  1140
lead wall  lead  550  420  15  980

Component  Material  r_{in}  r_{out}  Length Distance from detector
(shape: cylinderical)  (cm)  (cm)  (cm)  assembly (cm)
primary collimator  tungsten  0.25  10  30  2080
secondary collimator  nickel  0.50  10  50  1500
vacuum pipe  iron  2.2  2.5  1438  na

TABLE III: Dimensions of the light shielded collimator components.

Component  Material Width Height Length Distance from detector
(shape: cubic)  (cm)  (cm)  (cm)  assembly (cm)
concrete housing  concrete  550  420  1200  na
inner collimator cave  air  450  270  1150  na
first sweeping magnet  iron  20  30  50  2020
first absorber  iron  50  50  200  1810
second sweeping magnet  iron  20  30  50  1155
second absorber  iron  100  100  100  1045
lead wall  lead  450  270  15  1030

Component  Material  r_{in}  r_{out}  Length Distance from detector
(shape: cylinderical)  (cm)  (cm)  (cm)  assembly (cm)
primary collimator  tungsten  0.25  10  30  2080
secondary collimator  nickel  0.50  10  50  1215
vacuum pipe  iron  2.2  2.5  1203  na

TABLE IV: Background rates for the heavy shielding option.

particle type  background rate  background occupation  total background
in barrel DC (s^{−1})  fraction in barrel DC  through DET7 (s^{−1})
photons  (9.52±0.12)×10^{5}  0.133±0.002  (1.05±0.003)×10^{7}
e^{−} e^{+}  (8.03±0.34)×10^{4}  (1.20±0.05)×10^{−2}  (5.02±0.09)×10^{5}
neutrons  (1.19±0.13)×10^{4}  (2.0±0.2)×10^{−3}  (8.60±0.35)×10^{4}
μ^{−} μ^{+}  (2.60±0.61)×10^{3}  (4.0±0.9)×10^{−4}  (1.33±0.14)×10^{4}
total  (1.05±0.01)×10^{6}  (0.145±0.002)  (1.11±0.004)×10^{7}

FIG. 6: Graphs of the flux near the detector assembly for various particles with the heavy shielding option

TABLE V: Background rates for the moderate shielding option.

particle type  background rate  background occupation  total background
in barrel DC (s^{−1})  fraction in barrel DC  through DET7 (s^{−1})
photons  (9.07±0.11)×10^{5}  0.127±0.002  (1.09±0.004)×10^{7}
e^{−} e^{+}  (7.30±0.32)×10^{4}  (1.1±0.05)×10^{−2}  (5.91±0.09)×10^{5}
neutrons  (9.00±1.13)×10^{3}  (1.35±0.17)×10^{−3}  (6.70±0.31)×10^{4}
μ^{−} μ^{+}  (4.43±0.80)×10^{3}  (6.64±1.19)×10^{−4}  (1.53±0.15)×10^{4}
total  (9.93±0.12)×10^{5}  0.138±0.002  (1.16±0.004)×10^{7}

FIG. 7: Graphs of flux near the detector assembly for various particles with the moderate shielding option

TABLE VI: Background rates for the light shielding option.

particle type  background rate  background occupation  total background
in barrel DC (s^{−1})  fraction in barrel DC  through DET7 (s^{−1})
photons  (1.47±0.005)×10^{7}  0.890±0.003  (3.15±0.002)×10^{8}
e^{−} e^{+}  (8.45±0.11)×10^{5}  0.119±0.002  (1.93±0.005)×10^{7}
neutrons  (1.14±0.13)×10^{4}  (1.71±0.19)×10^{−3}  (1.01±0.04)×10^{5}
μ^{−} μ^{+}  (4.57±0.81)×10^{3}  (7.00±1.24)×10^{−4}  (2.60±0.19)×10^{4}
total  (1.56±0.005)×10^{7}  (0.903±0.003)  (3.34±0.002)×10^{8}

FIG. 8: Graphs of flux near the detector assembly for various particles with the light shielding option